from the top of a tower 40m high a ball is projected upward with speed 20m/s at an angle of elevation of 30 degree . calculate the time taken by the ball to hit the ground
 

Grade:11

1 Answers

Vikas TU
7 years ago
Write the eqns. for y- direction,
-40 = 20sin30*t – 0.5gt^2
-40 = 10*t – 5t^2
t^2 – 2t – 8 = 0
t^2 -4t + 2t – 8 = 0
t(t-4) + 2(t-4) = 0
(t+2) (t-4) = 0
t = 4 seconds.